# First Heuristic Then Rational: Dynamic Use of Heuristics in Language Model Reasoning

### v1

@@ Results:
The results are shown in Figure 3. The x-axis is the remaining steps $d$ to the goal, and the y-axis is the ratio $r$. The more distant the current step is from the answer (larger $d$), the more frequently the distractor is selected (larger $r$) which is typically above the chance rate. PaLM2 and GPT-4 exhibited particularly clear tendencies of the negative slopes between $d$ and $r$. These suggest the systematic behavior that LMs tend to rely more on heuristics in the earlier reasoning steps.
@@ 5 Related work
Given the historical goal of neuro-symbolic integration Hamilton et al. (2022) and the increasing relevance of neural LMs Fang et al. (2024), whether neural models can emulate particular symbolic operations (e.g., search over the graph) has been a key question Yao et al. (2023). In contrast to delineating what symbolic tasks are (im)possible for LMs by varying task complexities Clark et al. (2020), we investigate the inherent, systematic biases in solving a certain reasoning task.
@@
Neural models have typically been distracted by superficial biases Du et al. (2022). For example, they tend to use superficial features such as overlaps Lai et al. (2021); Sen and Saffari (2020), positional Ko et al. (2020), and specific syntactic category features Du et al. (2021); Niven and Kao (2019) even with chain-of-thought prompting Madaan and Yazdanbakhsh (2022); these motivated our experimental settings.
@@
Finding the shortest path between the start and the goal on a graph is a standard problem in computer science Russell and Norvig (2016). Our investigation of LMs on the arithmetic tasks can be seen as characterizing LMs’ biases as a search algorithm. The use of heuristics in graph search is, more or less, related to the A* search algorithm Hart et al. (1968), although heuristics in A* search is a more narrow concept regarding the distance to the goal than those employed by LMs, e.g., position bias.
@@ 6 Conclusion
We have found a systematic strategy for the use of heuristics in LMs’ multi-step reasoning—a dynamic transition from a heuristic to a rational reasoning strategy during the course of LMs’ step-by-step reasoning. These results are hopefully helpful for researchers to understand their underlying mechanism as well as for LM users to consider the inherent biases systems have.
@@
This study focused only on four specific language models and two arithmetic tasks. Increasing the coverage of models and tasks are obviously possible future direction, although we ensured that our finding generalizes at least several models and task settings. Nevertheless, in particular, Section 4.2 used only artificial datasets for fair comparison. Constructing a controlled, but natural dataset to evaluate the reasoning strategies of LMs should be encouraged. Furthermore, our findings are based solely on the model’s output texts. Elucidating the underlying mechanisms and the source of these behaviors (e.g., training data) should be investigated in the future work.
@@
This paper does not involve ethical concerns in the sense that we (i) did not conduct human experiments, (ii) just created artificial data without any potentially harmful contents, and (iii) did not address tasks related to ethically sensitive topics (i.e., arithmetic reasoning).
@@ Acknowledgements
We thank for the member of Tohoku NLP Group for their cooperation in this research.
This work was supported by JSPS KAKENHI Grant Numbers 21K21343 and 24K16077, JST CREST Grant Number JPMJCR20D2 and JST SPRING Grant Number JPMJSP2114.
Appendix A Experimental setting with GSM8K (§ 4.1)
@@ A.1 Dataset construction process
As described in 4.1, we modify the existing multi-hop numerical reasoning dataset, GSM8K (distributed under the MIT license), to construct the evaluation dataset. The dataset construction process is divided into two steps: 1. Extracting samples that can be used as evaluation data for this study, and 2. Inserting distractors according to the heuristic we want to evaluate for each extracted problem statement.
A.1.1 Sample extraction
@@
We extract samples from the GSM8K evaluation dataset for this study following the process below:
@@
    Using regular expressions, we identify PNs from this list, as well as pronouns and numerical expressions present in each sample.
@@
    We extracted samples that included exactly one from our list in both the context and the question, and where either the PN or a pronoun appears in the question (e.g., Table 3).
@@
    We replaced all pronouns within the extracted samples with the corresponding person’s name.
@@ A.1.2 Distractor insertion
Subsequently, we added distractors to the extracted samples according to each heuristic, thereby constructing a total of 76 samples for the evaluation dataset. Below, we will describe the process of creating the evaluation dataset for each heuristic.
@@
As a baseline, we insert a template-based random distractor (i.e., $\tilde{p}$) into each sample. The distractor was created using the following steps:
@@
    We randomly selected one sentence from the sample that included both a PN or pronoun and a numerical expression.
@@
    We replaced [name]  with a randomly selected name from the list of PNs created in A.1.1, excluding the name already present in the sample.
@@
    We inserted the created distractor into a random position in the context other than the beginning of the sample.
For example, When the sentence “James decides to run 3 sprints 3 times a week.” is selected from the sample in Table 3, a template “[name]  decides to run [num]  sprints [num]  times a week.” is crafted. Names and numbers are randomly selected from the candidates and placed into these placeholders, and the resulting distractor is then inserted into the context.
@@
To evaluate whether the model is influenced by the Overlap heuristic, we insert distractors $\tilde{p}$ into each sample following the steps below:
@@
    We substituted the placeholder [name]  within the Base distractor template with the person’s name found in the sample, appended by relational phrases such as “’s mother”, “’s father”, “’s son”, or “’s neighborhood” (e.g., in the sample from Table 3, this would become “James’s mother”).
@@
    We placed the constructed distractor into context at the exact location where the Base distractor was positioned in the sample.
@@
To evaluate whether the model is induced by the heuristic of Position, we insert distractors $\tilde{p}$ into each sample. Each distractor is identical to the Base distractor except for its insertion point. Specifically, we relocated the insertion point of the distractor to a random position that is closer to the beginning of the context than the position used for the Base distractor.
Neg.
To evaluate the model’s response to the Neg. heuristic, we insert distractors $\tilde{p}$ into each sample created based on the following template:
@@
  • [name]  is substituted with a random PN included in the sample.
@@
  • [num]  is replaced with a different numerical value, using the same algorithm used for creating the Base distractor.
@@ A.2 Evaluation
To determine whether the LMs selected the distractor during reasoning, we check if the numbers contained in the distractor $\tilde{p}$ are present in the facts $\bm{z}$. We calculate the percentage of samples where the distractor is selected.
Appendix B Experimental setting with artificial data (§ 4.1)
@@ B.1 Dataset construction process
We construct the artificial data by following the method outlined below, based on the template presented in Table 4.
@@
  • Assign either “more” or “less” to the placeholder [relation].
@@
We constructed a dataset to evaluate whether the model is induced by the Overlap heuristic by making certain modifications to the Base distractor for each sample. Specifically, we modified the value of [nameD]  by appending relational phrases such as “ ’s mother”, “ ’s father”, “ ’s son”, or “ ’s neighborhood” to the existing value of [nameD]. We then assigned this modified value to [nameE].
@@
To evaluate if the model is induced by the Position heuristic, we modify the Base distractor. Specifically, we altered the insertion point of the Base distractor to a randomly chosen position that is closer to the beginning of the context than the original position used in the Base distractor.
Neg.
To evaluate whether the model is induced by the Neg. heuristic, we construct a dataset by making modifications to the Base distractor. Specifically, we convert the Base distractor into a negative expression (e.g., [nameE]  doesn’t have [num]  [relation]  [object]  than [nameX]).
@@ B.2 Evaluation
To determine whether the LMs selected the distractor during reasoning, we check if the subject of the distractor (i.e., [nameE]) is included in the facts $\bm{z}$. We calculate the percentage of samples where the distractor is selected.
@@
  • The placeholder [relation]  is assigned either “more” or “less”.
@@
Then, using the following procedures, We create each expanded dataset. The heuristic distractors designed in this study are strongly influenced by each targeted heuristic. Each dataset consists of 300 problems.
@@
  • For the Neg. dataset, the form of the heuristic distractor is changed to a negative form.
@@
In the experimental section, we not only measure the frequency with which $\tilde{p}$ is used in reasoning but also evaluate the accuracy. Table 6 below shows the accuracy rates of each model on GSM8K. Additionally, Table 7 below shows the accuracy rates of each model on artificial data.
From the Table 67, it is shown that GPT-4 had the highest accuracy rates across the datasets, while Llama2 had the lowest. It is expected that these outcomes are due to differences in the number of parameters in the model and the training data.
@@ Appendix E Generation settings
When using GPT-3.5, GPT-4, the settings are adjusted to temperature=0.0, frequency_penalty=0, and presence_penalty=0. Similarly, for PaLM2 and Llama2, the temperature is set to 0, with no sampling.
@@ Appendix F Few-shot examples
Answer the context question according to the following example.
@@
Context: Betty is saving money for a new wallet which costs \$100. Betty has only half of the money she needs. Alice is saving money for a new wallet which costs \$2000. Bettys parents decided to give her \$15 for that purpose, and her grandparents twice as much as her parents. Question: How much more money does Betty need to buy the wallet?
@@
This means, Betty needs 100 - 50 - 30 - 15 = 5 more.
@@
Context: Julie is reading a 120-page book. Yesterday, Julie was able to read 12 pages and today, she read twice as many pages as yesterday. Julie’ s mother makes \$18.00 an hour.
Question: How many pages are left to be read?
@@
So she was able to read a total of 12 + 24 = 36 pages since yesterday.
@@
Since she wants to read half of the remaining pages tomorrow, then she should read 84/2 = 42 pages.
@@
Context: James writes a 2-page letter to 4 different friends, lived in America, twice a week. James writes a 3-page letter to 2 different friends, lived in Japan, twice a week.
Question: How many pages does James write each friend lived in Japan at one time?
@@
So he writes 6*2=12 pages every week.
@@
Answer the context question according to the following example.
@@
Context: Alice has 92 more bananas than Mallory. Victor has 10 less bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne has 79 more bananas than Sybil. Judy has 23 more bananas than Alice. Dave has 60 more bananas than Victor. Quentin has 35 less bananas than Peggy. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Peggy. Larry has 17 less bananas than Alice. Zoe has 58 less bananas than Yvonne. Ivan has 43 less bananas than Yvonne. Walter has 43 less bananas than Mallory. Nancy has 34 bananas. Grace has 41 more bananas than Xavier. Mallory has 55 less bananas than Nancy. Sybil has 3 less bananas than Nancy. Peggy has 50 more bananas than Walter. Trent has 33 less bananas than Xavier.
@@
Nancy has 34 bananas, and Mallory has 55 less bananas than Nancy. So, Mallory has 34-55=-21 bananas.
Mallory has -21 bananas, and Walter has 43 less bananas than Mallory. So, Walter has -21-43=-64 bananas.
@@
Peggy has -14 bananas, and Quentin has 35 less bananas than Peggy. So, Quentin has -14-35=-49 bananas.
@@
Context: Zoe has 10 more apples than Yvonne’s son. Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Quentin has 3 more apples than Yvonne. Yvonne has 3 less apples than Zoe. Alice has 3 more apples than Grace. Trent has 34 more apples than Zoe. Ivan has 3 apples. Ursula has 3 more apples than Zoe. Grace has 3 apples. Xavier doesn’t have 3 more apples than Ivan.
@@
Zoe has 15 apples, and Yvonne has 3 less apples than Zoe. So, Yvonne has 15-3=12 apples.
@@
Context: Kevin’s friend has 33 less grapes than Rob. Ivan has 43 more grapes than Victor. Victor has 33 less grapes than Kevin’s friend. Ursula has 75 less grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia has 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy has 78 more grapes than Olivia. Rob has 55 grapes. Frank has 70 less grapes than Heidi. Eve has 84 less grapes than Sybil. Xavier has 36 more grapes than Heidi. Sybil has 55 less grapes than Trent. Kevin has 43 less grapes than Zoe. Heidi has 61 less grapes than Trent. Zoe has 88 more grapes than Sybil. Trent has 40 more grapes than Rob. Walter has 38 more grapes than Victor.
@@
Trent has 95 grapes, and Sybil has 55 less grapes than Trent. So, Sybil has 95-55=40 grapes.
@@
Zoe has 128 grapes, and Kevin has 43 less grapes than Zoe. So, Kevin has 128-43=85 grapes.
The final answer is 128.

### v2

@@ Results:
The results are shown in Figure 3. The x-axis is the change of remaining steps $d$ to the goal in the respective reasoning step, and the y-axis is the ratio $r$ of selecting distractors $\tilde{p}_{t}$. The more distant the current step is from the answer (larger $d$), the more frequently the distractor is selected (larger $r$), which is typically above the chance rate. PaLM2 and GPT-4 exhibited apparent tendencies of the negative slopes between $d$ and $r$. These results match the hypothesized behavior, the model’s less rational behavior in earlier reasoning steps, and imply that they have a limited capacity to track the future reasoning path.
@@ 5 Related work
Given the general contrast between the symbolic and neural-based approaches in the artificial intelligence field Hamilton et al. (2022), the community has questioned the ability of neural LMs in emulating particular symbolic operations, e.g., graph search algorithm Aoki et al. (2023); Yao et al. (2023); Fang et al. (2024). In contrast, to identify what kind of symbolic tasks are (im)possible to solve for LMs by varying task complexities Clark et al. (2020), we investigate the inherent, systematic biases in solving a particular symbolic reasoning task.
@@
Neural models have typically been distracted by superficial biases Du et al. (2022). For example, they tend to use superficial linguistic artifacts Lai et al. (2021); Sen and Saffari (2020); Du et al. (2021); Niven and Kao (2019), or more simply, positional features Ko et al. (2020), even with chain-of-thought prompting Madaan and Yazdanbakhsh (2022); these motivated our experiments. This paper shows that the LLMs’ reliance on such heuristics changes dynamically as the reasoning progresses.
@@
Finding the shortest path between the start and the goal on a graph is a standard problem in computer science Russell and Norvig (1995). Our investigation of LMs on the arithmetic tasks can be seen as characterizing LMs’ biases as a search algorithm. The use of heuristics in graph search is, more or less, related to the A* search algorithm Hart et al. (1968), although heuristics in A* search is a more narrow concept regarding the distance to the goal than those employed by LMs.
@@ 6 Conclusion
We have found a systematic strategy for the use of heuristics in LMs’ multi-step reasoning—a dynamic transition from a heuristic to a rational reasoning strategy during LMs’ step-by-step reasoning. These results are hopefully helpful for researchers to understand their underlying mechanism as well as for LM users to consider the inherent biases systems have.
@@
This study focused only on four specific language models and two arithmetic tasks. Increasing the coverage of models and tasks is a possible future direction, although we ensured that our finding generalizes at least several models and task settings. In § 4.2, we only used artificial datasets for designing controlled experiments to reduce confounding factors. Constructing a controlled but natural dataset to evaluate the reasoning strategies of LMs should be encouraged. Furthermore, our findings are based solely on the model’s outputs, i.e., behaviors. Elucidating the underlying mechanisms inside the model and the source of these biases (e.g., statistical patterns in training data) should be investigated in future work.
@@
This paper does not involve ethical concerns in the sense that we (i) did not conduct human experiments, (ii) just created artificial data without any potentially harmful contents, and (iii) did not address tasks related to ethically sensitive topics.
@@ Acknowledgements
We thank the member of the Tohoku NLP Group for their cooperation in this research. This work was supported by JSPS KAKENHI Grant Numbers 21K21343 and 24K16077, JST CREST Grant Number JPMJCR20D2, JST SPRING Grant Number JPMJSP2114, and JST BOOST Grant Number JPMJBS2421.
Appendix A GSM8K experiments in § 4.1
@@ A.1 Dataset construction process
As described in 4.1, we modify the existing multi-hop numerical reasoning dataset, GSM8K (distributed under the MIT license), to construct the evaluation dataset. The dataset construction process is divided into two steps: 1. Extracting instances from GSM8K, and 2. Inserting distractors according to the heuristic we want to evaluate for each extracted problem statement.
A.1.1 Instance extraction
@@
There is no guarantee that the premises added during data expansion will not affect the solution. Therefore, instances are extracted in which the addition of premises has little effect on the solution. Specifically, we extract instances from the GSM8K evaluation dataset following the process below:
@@
    Using regular expressions, we identify PNs from this list.
@@
    We identify pronouns and numerical expressions present in each instance.
@@
    We extracted instances that included precisely one from our list in both the context and the question, and where either the PN or a pronoun appears in the question (e.g., Table 3).
  5. 5.
    We replaced all pronouns within the extracted instances with the corresponding person’s name.
In the following process, information on persons not appearing in the problem statement is added as distractors. The instance questions extracted above are about persons in the problem statement. Therefore, it can be guaranteed that adding a distractor will not change the answer by such a process.
@@ A.1.2 Distractor insertion
Subsequently, we added distractors to the extracted instances according to each heuristic, thereby constructing 76 instances for the evaluation dataset. Below, we will describe the process of creating the evaluation dataset for each heuristic.
@@
We insert a template-based random distractor (i.e., $\tilde{p}$) into each instance as a baseline. The distractor was created using the following steps:
@@
    We randomly selected one sentence from the instance that included a PN or pronoun and a numerical expression.
@@
    We replaced [name]  with a randomly selected name from the list of PNs created in A.1.1, excluding the name already present in the instance.
@@
    We inserted the created distractor into a random position in the context other than the beginning of the instance.
For example, When the sentence “James decides to run 3 sprints 3 times a week.” is selected from the instance in Table 3, a template “[name]  decides to run [num]  sprints [num]  times a week.” is crafted. Names and numbers are randomly selected from the candidates and placed into these placeholders, and the resulting distractor is then inserted into the context.
@@
To evaluate whether the Overlap heuristic influences the model, we insert distractors $\tilde{p}$ into each instance following the steps below:
@@
    We substituted the placeholder [name]  within the Base distractor template with the person’s name found in the instance, appended by relational phrases such as “’s mother”, “’s father”, “’s son”, or “’s neighborhood” (e.g., in the instance from Table 3, this would become “James’s mother”).
@@
    We placed the constructed distractor into context at the location where the Base distractor was positioned in the instance.
@@
To evaluate whether the heuristic of Position induces the model, we insert distractors $\tilde{p}$ into each instance. Each distractor is identical to the Base distractor except for its insertion point. Specifically, we relocated the distractor’s insertion point to a random position closer to the beginning of the context than the position used for the Base distractor.
Negative
To evaluate the model’s response to the Negative heuristic, we insert distractors $\tilde{p}$ into each instance created based on the following template:
@@
  • [name]  is substituted with a random PN included in the instance.
@@
  • [num]  is replaced with a different numerical value, using the same algorithm for creating the Base distractor.
@@ A.2 Evaluation
To determine whether the LMs selected the distractor during reasoning, we check if the numbers in the distractor $\tilde{p}$ are in the facts $\bm{z}$. We calculate the frequency of instances where the distractor is selected.
Appendix B Artificial-data experiments in § 4.1
@@ B.1 Dataset construction process
We construct the artificial data using the method outlined below, based on the template presented in Table 4.
@@
  • Assign either “more” or “fewer” to the placeholder [relation].
@@
We constructed a dataset to evaluate whether the Overlap heuristic induced the model by making certain modifications to the Base distractor for each instance. Specifically, we modified the value of [nameD]  by appending relational phrases such as “ ’s mother”, “ ’s father”, “ ’s son”, or “ ’s neighborhood” to the existing value of [nameD]. We then assigned this modified value to [nameE].
@@
We modify the Base distractor to evaluate if the Position heuristic induces the model. Specifically, we altered the insertion point of the Base distractor to a randomly chosen position closer to the context’s beginning than the original position used in the Base distractor.
Negative
To evaluate whether the Neg induces the model heuristic, we construct a dataset by modifying the Base distractor. Specifically, we convert the Base distractor into a negative expression (e.g., [nameE]  doesn’t have [num]  [relation]  [object]  than [nameX]).
@@ B.2 Evaluation
To determine whether the LMs selected the distractor during reasoning, we check if the subject of the distractor (i.e., [nameE]) is included in the facts $\bm{z}$. We calculate the frequency of instances where the distractor is selected.
@@
  • The placeholder [relation]  is assigned either “more” or “fewer”.
@@
Then, using the following procedures, We create each expanded dataset. Each targeted heuristic strongly influences the heuristic distractors designed in this study. Each dataset consists of 300 problems.
@@
  • For the Negative dataset, the form of the heuristic distractor is changed to a negative form.
@@
This paper was mainly concerned with the frequency of distractor selection. To ensure that the model is not producing crappy output in these experiments, we measure the accuracy. Table 6 below shows the accuracy rates of each model on GSM8K. Additionally, Table 7 below shows the accuracy rates of each model on artificial data.
From the Table 67, it is shown that GPT-4 had the highest accuracy rates across the datasets, while Llama2 had the lowest. It is expected that these outcomes are due to differences in the number of parameters in the model.
@@ Appendix E Generation settings
When using GPT-3.5 and GPT-4, the settings are adjusted to temperature=0.0, frequency_penalty=0, and presence_penalty=0. Similarly, for PaLM2 and Llama2, the temperature is set to 0, with no sampling.
@@ Appendix F Few-shot examples
Answer the context question using the following example.
@@
Context: Betty is saving money for a new wallet, which costs \$100. Betty has only half of the money she needs. Alice is saving money for a new wallet, which costs \$2000. Betty’s parents decided to give Betty \$15 for that purpose, and her grandparents twice as much as her parents. Question: How much more money does Betty need to buy the wallet?
@@
This means Betty needs 100 - 50 - 15 - 30 = 5 more.
@@
Context: Julie is reading a 120-page book. Yesterday, Julie was able to read 12 pages, and today, she read twice as many pages as yesterday. Julie’s mother makes \$18.00 an hour.
Question: If Julie wants to read half of the remaining pages tomorrow, how many are left to read?
@@
So, she was able to read a total of 12 + 24 = 36 pages since yesterday.
@@
Since she wants to read half of the remaining pages tomorrow, she should read 84/2 = 42 pages.
@@
Context: James writes a 2-page letter to 4 different friends who live in America twice a week. James writes a 3-page letter to 2 different friends who live in Japan twice a week.
Question: How many pages does James write each friend who lived in Japan for a year?
@@
So, he writes 6*2=12 pages every week.
@@
Answer the context question using the following example.
@@
Context: Alice has 92 more bananas than Mallory. Victor has 10 fewer bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne has 79 more bananas than Sybil. Judy has 23 more bananas than Alice. Dave has 60 more bananas than Victor. Quentin has 35 fewer bananas than Peggy. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Peggy. Larry has 17 fewer bananas than Alice. Zoe has 58 fewer bananas than Yvonne. Ivan has 43 fewer bananas than Yvonne. Walter has 43 fewer bananas than Mallory. Nancy has 34 bananas. Grace has 41 more bananas than Xavier. Mallory has 55 fewer bananas than Nancy. Sybil has 3 fewer bananas than Nancy. Peggy has 50 more bananas than Walter. Trent has 33 fewer bananas than Xavier.
@@
Nancy has 34 bananas, and Mallory has 55 fewer bananas than Nancy. So, Mallory has 34-55=-21 bananas.
Mallory has -21 bananas, and Walter has 43 fewer bananas than Mallory. So, Walter has -21-43=-64 bananas.
@@
Peggy has -14 bananas, and Quentin has 35 fewer bananas than Peggy. So, Quentin has -14-35=-49 bananas.
@@
Context: Zoe has 10 more apples than Yvonne’s son. Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Quentin has 3 more apples than Yvonne. Yvonne has 3 fewer apples than Zoe. Alice has 3 more apples than Grace. Trent has 34 more apples than Zoe. Ivan has 3 apples. Ursula has 3 more apples than Zoe. Grace has 3 apples. Xavier doesn’t have 3 more apples than Ivan.
@@
Zoe has 15 apples, and Yvonne has 3 fewer apples than Zoe. So, Yvonne has 15-3=12 apples.
@@
Context: Kevin’s friend has 33 fewer grapes than Rob. Ivan has 43 more grapes than Victor. Victor has 33 fewer grapes than Kevin’s friend. Ursula has 75 fewer grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia has 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy has 78 more grapes than Olivia. Rob has 55 grapes. Frank has 70 fewer grapes than Heidi. Eve has 84 fewer grapes than Sybil. Xavier has 36 more grapes than Heidi. Sybil has 55 fewer grapes than Trent. Kevin has 43 fewer grapes than Zoe. Heidi has 61 fewer grapes than Trent. Zoe has 88 more grapes than Sybil. Trent has 40 more grapes than Rob. Walter has 38 more grapes than Victor.
@@
Trent has 95 grapes, and Sybil has 55 fewer grapes than Trent. So, Sybil has 95-55=40 grapes.
@@
Zoe has 128 grapes, and Kevin has 43 fewer grapes than Zoe. So, Kevin has 128-43=85 grapes.
The final answer is 85.